OpenStudy (eric_d):

Find range of values of x such that each of the following inequalities is valid

OpenStudy (eric_d):

\[x(5x^2+8) \le 1/2(47x^2-48)\]

OpenStudy (eric_d):

\[10x^3-47x^2+16x+48 \le 0\]

OpenStudy (eric_d):

What shall I do nxt @ganeshie8

ganeshie8 (ganeshie8):

I would try to factor it

OpenStudy (eric_d):

After that

OpenStudy (eric_d):

Nvm, I gt the answer

OpenStudy (eric_d):

Can we factorise the above quation without using a calculator

OpenStudy (eric_d):

equation*

ganeshie8 (ganeshie8):

yes but its a bit lengthy process, you need to know below goodies : 1) descartes rule of signs 2) rational root theorem 3) synthetic division

OpenStudy (eric_d):

I'll jst use ur first method

OpenStudy (eric_d):

What about this question 2x^2-3x-12 smaller or equal to 0 Same method as above to find the inequalities?

ganeshie8 (ganeshie8):

may be factor it by using quadratic formula

OpenStudy (eric_d):

The values might be in decimals But the answer given is \[x \le -3/2,2<x \le4\]

ganeshie8 (ganeshie8):

check the equation once.. is the equation really a quadratic ?

OpenStudy (eric_d):

yes..

ganeshie8 (ganeshie8):

not sure, im also getting decimals...